Evaluate the limit as x turns to -∞

Lim [3x + √(9x2-Ix-2I)]


1
Expert's answer
2021-10-28T07:09:14-0400
"\lim\limits_{x\to-\infin}(3x+\sqrt{9x^2-|x-2|})"

"=\lim\limits_{x\to-\infin}(3x+\sqrt{9x^2-|x-2|})\cdot\dfrac{3x-\sqrt{9x^2-|x-2|}}{3x-\sqrt{9x^2-|x-2|}}"

"=\lim\limits_{x\to-\infin}\dfrac{9x^2-9x^2+|x-2|}{3x-\sqrt{9x^2-|x-2|}}"

"=\lim\limits_{x\to-\infin}\dfrac{\dfrac{|x-2|}{|x|}}{\dfrac{3x}{|x|}-\sqrt{\dfrac{9x^2}{x^2}-\dfrac{|x-2|}{x^2}}}"

"=\dfrac{1}{-3-\sqrt{9-0}}=-\dfrac{1}{6}"